OpenStudy (anonymous):

Please help ^_^ Questions are in the comments. Thank you!!! 1) 14 3/8 gal.= _____ qt. 2) 15 3/4 pt.= _____ qt. 3) 3 gal.= ____ fl. oz. 4) 8 1/2 pt.= ____ fl. oz. 5) 19.3 kL= _____ hL 6) 3 L= _____ mL

OpenStudy (anonymous):

1) 14 3/8 gal.= _____ qt. 2) 15 3/4 pt.= _____ qt. 3) 3 gal.= ____ fl. oz. 4) 8 1/2 pt.= ____ fl. oz. 5) 19.3 kL= _____ hL 6) 3 L= _____ mL

OpenStudy (mertsj):